I need help on knowing how to solve linear systems in form of fractions.
the problem: (1/x+2) + (2/y+1) = 1
(4/x+2) + (9/y+1) = 13/3
i attempted the first equation
i multiplied the left side by (x+2) (y+1) and the right side as well.
so i got y+1+2x+4=xy+x+2y+2
then i subtracted x and y from both sides
5+x=xy+y+2
to get
xy+y-x-3=o
I do not know if this is correct, so i can't go on solving the problem because i don't know. Truthfully, im guessing at this point.

Found 2 solutions by user_dude2008, amoresroy:
Answer by user_dude2008(1862)

(1/x+2) + (2/y+1) = 1 -----> y+1+2(x+2)=(x+2)(y+1)


(4/x+2) + (9/y+1) = 13/3 ----> 12(y+1)+27(x+2)=13(x+2)(y+1)

Answer by amoresroy(361)

First Equation
Multiplying both sides of equation by (x+2) (y+1), gives you
1 (y+1) + 2 (x+2) = (x+2) (y+1)
y+1+2x+4 = xy + x + 2y + 2
xy - x +y = 3
So far your solution is correct. Proceed to simply the second equation.